OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 1,024 cubic inches and has a height of 12 inches. Use 3.14 for pi and round your answer to the nearest hundredth. Select one: a. 4 inches b. 8 inches c. 16 inches d. 24 inches

OpenStudy (lastdaywork):

See the formula for volume of (general) square pyramid - http://en.wikipedia.org/wiki/Square_pyramid#Other_square_pyramids Can you answer the question now ??

OpenStudy (anonymous):

What equation do I use? @LastDayWork

OpenStudy (lastdaywork):

For square pyramids in general, with base length l and height h, volume is: \[V=\frac{ 1 }{ 3 } l^2 h\]

OpenStudy (anonymous):

So then would it be set up like: 1024=1/3L(12)

OpenStudy (anonymous):

@LastDayWork

OpenStudy (lastdaywork):

Yep..the equation is correct. We can now solve for l :)

OpenStudy (anonymous):

Okay...So first I would multiply 1/3 and 12, which comes out to 4. Then the equation is 1024=4L. Now what?

OpenStudy (anonymous):

@LastDayWork

OpenStudy (lastdaywork):

A correction; it's L^2 So \[1024 = 4L^2\] We can add or subtract a number on both sides of equal to. We can multiply or divide a (non-zero) number on both sides of equal to So, \[1024 = 4L^2\] \[\frac{ 1024 }{ 4 }=\frac{ 4L^2 }{ 4 }\] Can you simplify it ??

OpenStudy (anonymous):

It would be 16 :) Because 1024/4 is 256, and the square root of that is 16 :)

OpenStudy (lastdaywork):

That's correct :)